OpenStudy (anonymous):

a bakery charges .10 for a cookie that is 2 inches in diameter. if the price is proportional to the area how much do they charge for an extra large cookie that is 8 inches in diameter? Assume that each cookie is shaped of a circle a..40 b.80 c. 1.60 d. 6.40 if you answer it you get a medal

OpenStudy (anonymous):

Start by finding the area of the two cookies. The area of a circle is equal to pi times the square of the radius: \[A=\pi r^2\]So what do you get for each of the areas? Remember that the radius is one half of the diameter.

OpenStudy (anonymous):

idk @vinnv226

OpenStudy (tkhunny):

Or, directly \(\dfrac{\pi\cdot 4^{2}}{\pi\cdot 1^{2}} = 16\)

OpenStudy (texaschic101):

I am not sure if this is correct, but it says proportional....I am going to make a proportion. 10/2 = x/8 ($10 to 2 in = $x to 8 inches) cross multiply (2)(x) = (10)(8) 2x = 80 x = 40 Like I said, I am not sure if this is correct, so you will need another opinion.

OpenStudy (tkhunny):

No, @texaschic101, the proportion is in the AREA, not he Diameter.

OpenStudy (texaschic101):

sorry.....I tried :(